Find the intensity of an 89 dB sound.  What happens to this intensity if you move 4 times the distance away (actually need a value)? What is the dB reading of this new intensity?



1
Expert's answer
2021-06-15T10:02:16-0400

1)


"89=10\lg\frac{I}{10^{-12}}\\I=0.000794\frac{W}{m^2}"

2)


"I'=\frac{1}{4^2}I=\frac{1}{4^2}0.000794=0.000049625\frac{W}{m^2}"

"\beta'=10\lg\frac{0.000049625}{10^{-12}}=77\ dB"